OpenStudy (anonymous):

whats the 21st term in the arithemtic sequence in which...

OpenStudy (anonymous):

\[a _{3} is 59 and a _{1} is 103 \]

OpenStudy (anonymous):

enter the term

OpenStudy (anonymous):

What is 59-103?

OpenStudy (anonymous):

term 1 = 103 term 2 = ? term 3 = 59 In order to find term two find the difference between term 1 and term 3 and divide that by 2. That will give you your change in value. then you will have everything you need to make another equation in the form of : An = term 1 +(change in value)*(n-1)

OpenStudy (anonymous):

once you are there plug in 21 for n and you will have your corresponding 21st term

OpenStudy (anonymous):

so -44 /2 = -22

OpenStudy (anonymous):

-463?

OpenStudy (anonymous):

correct! so that's you change in value. now you can make your formula

OpenStudy (anonymous):

but its asking for the term

OpenStudy (anonymous):

I got -337

OpenStudy (anonymous):

A_21 = 103-22*(21-1) = 103 - 440 = -337

OpenStudy (anonymous):

so the term is -337

OpenStudy (anonymous):

?

OpenStudy (anonymous):

You tell me

OpenStudy (anonymous):

yes?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

so -337?(:

OpenStudy (anonymous):

ahha you really need your verification don't you .. yes yes it is

OpenStudy (anonymous):

Yes(: